HIGH SENSITIVITY CRDS OF THE a^{1}Δ_{g}←X^{3}Σ_{g}^{−} BAND OF OXYGEN NEAR 1.27 μm: MAGNETIC DIPOLE AND ELECTRIC QUADRUPOLE TRANSITIONS IN SPECTRA OF FIVE ISOTOPOLOGUES

O. M. LESHCHISHINA, S. KASSI, L. WANG, Université Joseph Fourier/CNRS, Laboratoire de Spectrométrie Physique, 38402 Saint Martin d’Hères, FRANCE; \underline{I. E. GORDON}, L. S. ROTHMAN, Harvard-Smithsonian Center for Astrophysics, Atomic and Molecular Physics Division, Cambridge MA 02138-1516, USA; A. CAMPARGUE, Université Joseph Fourier/CNRS, Laboratoire de Spectrométrie Physique, 38402 Saint Martin d’Hères, FRANCE.

The knowledge of accurate spectroscopic parameters for the a^{1}Δ_{g}←X^{3}Σ_{g}^{−} band of molecular oxygen near 1.27 μm is very important in the field of remote sensing. Although this band was studied by spectroscopists for over a century a lot of discrepancies still remain in the previously reported line positions and intensities. In this work the Continuous Wave-Cavity Ring Down Spectroscopy (CW-CRDS) technique has been used to record with high sensitivity the absorption spectrum of this band. The spectra were obtained between 7640 and 7917 cm^{−1} with “natural” oxygen and with a sample highly enriched in ^{18}O. The absolute intensities of 377 and 652 oxygen transitions were measured in the two spectra, respectively. They include the a^{1}Δ_{g}←X^{3}Σ_{g}^{−} (0-0) bands of ^{16}O_{2}, ^{16}O^{18}O, ^{16}O^{17}O, ^{17}O^{18}O and ^{18}O_{2}. The (0-0) bands of ^{16}O_{2} and ^{18}O_{2} show (previously undetected) electric quadrupole transitions with line intensities ranging from 1×10^{−30} to 1.9×10^{−28} cm/molecule. They are accompanied by the a^{1}Δ_{g}←X^{3}Σ_{g}^{−} (1-1) hot bands which are also reported for the first time. Accurate spectroscopic parameters for the observed bands were derived from a global fit of the experimental line positions, combined with microwave and Raman measurements available in the literature.